# Parameter estimation for Gibbs distributions

### v9

@@ 2.1 The Balance subroutine
Given a target $\chi$, we need to find a value $\beta$ with $\mu_{\beta}[0,\chi]\approx 1/2\approx\mu_{\beta}[\chi,n]$. That is, $\chi$ is approximately the median of the distribution $\mu_{\beta}$. We can find the target value $\beta$ via noisy binary search.
Formally, for $\tau\in(0,1/2),\chi\in[0,n]$, we define ${\tt BalancingVals}(\chi,\tau)$ to be the set of values $\beta\in[\beta_{\min},\beta_{\max}]$ satisfying the following two properties:
  • Either $\beta=\beta_{\min}$ or $\mu_{\beta}[0,\chi]\geq\tau$
@@
Consider $\alpha\in[\beta_{\min},\beta_{\max}]$ with $\mu_{\alpha}(x)=\Delta(x)$. The result is clear if $\alpha=\beta$. Suppose that $\alpha<\beta$; the case $\alpha>\beta$ is completely analogous. So $\beta>\beta_{\min}$, and since $\beta\in{\tt BalancingVals}(x,\tau)$, this implies $\mu_{\beta}[0,x]\geq\tau$. We then have:
@@
In Appendix D, we show the following result:
Theorem 11.
Let $\tau\in(0,1/2)$ be an arbitrary constant. Then subroutine ${\tt Balance}(\chi,\tau,\gamma)$ has cost $O(\log\frac{nq}{\gamma})$. With probability at least $1-\gamma$, it returns a value $\beta\in{\tt BalancingVals}(\chi,\tau)$ (we say in this case that the call is good).
Theorem 12.
Let $\tau\in(0,1/2)$ be an arbitrary constant and let $\varepsilon=(1/2-\tau)/30$. If we are given a data structure $\mathcal{D}$ for ${P^{{\varepsilon}}_{\tt ratio-all}}$, then subroutine ${\tt BalancePreprocessed}(\chi,\tau,\mathcal{D},\gamma)$ has cost $O(\log\frac{q}{\gamma})$. If $\mathcal{D}$ is $\varepsilon$-close, then with probability at least $1-\gamma$, it returns a value $\beta\in{\tt BalancingVals}(\chi,\tau)$.
@@ 2.2 Statistical sampling
[…]
@@
Many of our algorithms are based on calling $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta,N)$ and making decisions depending on the values $\hat{\mu}_{\beta}(I)$ for certain sets $I\subseteq\mathcal{F}$; they succeed when the estimates $\hat{\mu}_{\beta}(I)$ are close to $\mu_{\beta}(I)$. We say the execution of Sample well-estimates $I$ if Eq. (2) hold for $p=\mu_{\beta}(I)$ and $\hat{p}=\hat{\mu}_{\beta}(I)$; otherwise it mis-estimates $I$. Likewise we say Sample well-estimates $k$ if it well-estimates the singleton set $I=\{k\}$. Since this comes up so frequently, we write
$\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta;\varepsilon,\bar{p},\gamma)$
as shorthand for $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta,\big{\lceil}\frac{3e^{\varepsilon}\log(4/\gamma)}{(1-e^{-\varepsilon})^{2}\bar{p}}\big{\rceil})$. It has cost $O(\frac{\log(1/\gamma)}{\varepsilon^{2}\bar{p}})$, and each individual set $I$ is well-estimated with probability at least $1-\gamma$. Here $\varepsilon$ should be interepreted as the desired relative error, $\bar{p}$ as the minimum “detection threshold” for $\mu_{\beta}$, and $\gamma$ as the desired failure probability.
In a similar way, our algorithms for ${P^{\delta,\varepsilon}_{\tt count}}$ estimate $\pi(x)$ by sampling $\hat{\mu}_{\alpha}(x)$ for a well-chosen value $\alpha$. We record the following straightforward error bound:
@@
Suppose that for $x\in\mathcal{F}$, we have $\alpha\in[\beta_{\min},\beta_{\max}]$ and non-negative parameters $\hat{Q}(\alpha),\hat{\mu}_{\alpha}(x)$ (which may depend upon $x$), satisfying the following bounds:
@@
Suppose we can draw independent samples from non-negative random variables $X_{1},\dots,X_{N}$. Given input $\gamma\in(0,1/2)$ and $\tau\geq 1$, the procedure ${\tt EstimateProducts}(X,\tau,\gamma)$ returns a vector of estimates $(\hat{X}^{\tt prod}_{1},\dots,\hat{X}^{\tt prod}_{N})$ using $O(N\tau\log\tfrac{1}{\gamma})$ total samples of the $X$ variables. If $\tau\geq\sum_{i=1}^{N}{{\mathbb{S}}[{X_{i}}]}$, then with probability at least $1-\gamma$, it holds that
for all $\text{for all $i=1,\dots,N$: }\frac{\hat{X}_{i}^{\tt prod}}{\prod_{j=1}^{i}{\mathbb{E}}[X_{j}]}\in[e^{-\varepsilon},e^{\varepsilon}]\qquad\qquad\text{for $\varepsilon=\sqrt{\sum_{i=1}^{N}{{\mathbb{S}}[{X_{i}}]}/\tau}\in(0,1)$}$
@@ 3 Solving ${P^{{\varepsilon}}_{\tt ratio-all}}$ in the continuous setting
3       sample $K\sim\mu_{\beta_{i-1}}$ and sample $U\sim\text{Unif}([0,1])$.
4       set $\beta_{i}=\beta_{i-1}+\frac{\log U}{K}$
5       if $K=0$ or $\beta_{i}<\beta_{\min}$ then return set $B=\{\beta_{1},\dots,\beta_{i-1}\}$
@@
3       draw $\eta$ from the Exponential distribution of rate $k$, and set $z_{i}=z_{i-1}-\eta$
@@
Let $\mathcal{E}$ denote the event that $\mathcal{D}$ fails to be good for some value $\alpha$. Consider the random process where $B\cap[\beta_{\min},\alpha]$ is revealed while $\alpha$ gradually increases; suppose $\mathcal{E}$ occurs and let $\alpha$ be the first value with $|B(\beta_{\min},\alpha)/k-z(\beta_{\min},\alpha)|>\varepsilon$. Since $z(B)$ is a Poisson process, $B\cap(\alpha,\beta_{\max})$ retains its original, unconditioned probability distribution. Accordingly, there is a probability of at least $21-\gamma/$ 2 that $|B(\alpha,\beta_{\max})/k-z(\alpha,\beta_{\max})|\leq\varepsilon/2$, in which case we would get
@@ 3.1 Paired Product Estimator
To improve the sample complexity to $O(\frac{q\log n}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$, we need a more advanced statistical technique called the Paired Product Estimator developed in [14]. Here, we describe a simple version of the algorithm. [^2] This is a simplified version of the PPE algorithm of [14]. See also [14, 22] for a different, stand-alone version of the PPE algorithm to directly solve ${P^{{\varepsilon}}_{\tt ratio-pt}}$ with high probability.
@@
2 define $\beta_{0}=\beta_{\min}$ and $\beta_{t}=\beta_{\min}$.
@@
4       define random variable $W_{i}=\exp\bigl{(}\frac{\beta_{i}-\beta_{i-1}}{2}\cdot K\bigr{)}$ where $K$ is drawn from $\mu_{\beta_{i-1}}$
5       define random variable $V_{i}=\exp\bigl{(}-\frac{\beta_{i}-\beta_{i-1}}{2}\cdot K\bigr{)}$ where $K$ is drawn from $\mu_{\beta_{i}}$
@@
There holds ${\mathbb{E}}[\kappa]\leq\frac{\log(z^{\prime}(\beta_{\max}))-\log(z^{\prime}(\beta_{\min}))}{k}.$
@@
As shown in [14, Lemma 3.2], we have $\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\geq\exp(\frac{2\kappa_{i}}{z(\beta_{i-1},\beta_{i})})$ for each $i$ (see also [25]). (Here $z^{\prime}(\beta)$ denotes the derivative of function $z$ at value $\beta$). Equivalently, we have $\kappa_{i}\leq\frac{z(\beta_{i-1},\beta_{i})}{2}\cdot\log\bigl{(}\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\bigr{)}$.
Observe that $z(\beta)$ and $z^{\prime}(\beta)$ are both strictly increasing functions of $\beta$. Thus, for any fixed value $x\in[z^{\prime}(\beta_{\min}),z^{\prime}(\beta_{\max})]$, there is a unique value $\beta\in[\beta_{\min},\beta_{\max}]$ with $z^{\prime}(\beta)=x$, and a unique index $i$ with $z(\beta)\in[z(\beta_{i-1}),z(\beta_{i}))$. Define $A_{x}=z(\beta_{i-1},\beta_{i})$, that is, the distance between the nearest points of the cooling schedule. We can rewrite the bound for $\kappa_{i}$ then as
@@
If $k\geq\frac{10(\log z^{\prime}(\beta_{\max})-\log z^{\prime}(\beta_{\min}))}{\varepsilon^{2}}$ for $\varepsilon\in(0,1)$, then with probability at least $0.7$, the following two bounds hold: (a) $\kappa\leq\varepsilon^{2}$ and (b) $|\log(\hat{W}_{i}^{\tt prod}/\hat{V}_{i}^{\tt prod})-\log Q(\beta_{i})|\leq\varepsilon/2$ for all $i=0,\dots,t$.
@@
$\frac{\hat{V}^{\tt prod}_{i}}{\prod_{j=1}^{i}{\mathbb{E}}[V_{j}]}\in[e^{-\varepsilon^{\prime},\varepsilon^{\prime}}],\qquad\frac{\hat{W}^{\tt prod}_{i}}{\prod_{j=1}^{i}{\mathbb{E}}[W_{j}]}\in[e^{-\varepsilon^{\prime},\varepsilon^{\prime}}]\qquad\text{for all $i=0,\dots,t$}$
where $\varepsilon^{\prime}=\varepsilon^{2}/4$. In this case, each value $\hat{Q}(\beta_{i})=\hat{W}^{\tt prod}_{i}/\hat{V}^{\tt prod}_{i}$ is within $[e^{-\varepsilon/2},e^{\varepsilon/2}]$ of the product $\prod_{j=1}^{i}\frac{{\mathbb{E}}[W_{j}]}{{\mathbb{E}}[V_{j}]}$; by telescoping products, this is precisely $Z(\beta_{i})/Z(\beta_{0})=Q(\beta_{i})$. ∎
@@
Consider $\alpha=(1-v)\beta_{i-1}+v\beta_{i}$ for $v\in[0,1]$. Suppose that $v\leq 1/2$; the case where $v\geq 1/2$ is completely symmetric. We need to show:
@@ 3.2 A hybrid algorithm
6       call $\mathcal{D}_{0}={\tt TPA-Pratio}(k_{1})$ with $k_{1}=\lceil\frac{40\cdot\log 4}{\varepsilon^{2}}\log\tfrac{12}{\gamma}\rceil$ for the interval $[\beta_{\min},\beta_{\circ}]$.
@@
8             Set $\mathcal{D}_{i}\leftarrow{\tt PPE-Pratio}(k_{2},\varepsilon/2)$ with $k_{2}=\lceil\frac{10\log(4n)}{\varepsilon^{2}}\rceil$ for the interval $[\beta_{\circ},\beta_{\max}]$
@@
output tuple $\mathcal{D}=(\beta_{\circ},\mathcal{D}_{1},\dots,\mathcal{D}_{t})$     Query algorithm $Q(\alpha\mid{\cal D})$:
@@
By Eq. (6), Theorem 16 and our choice of $k_{1}\geq\frac{80\log 4}{(\varepsilon/2)^{2}}\log\tfrac{3}{\gamma/4}$, with probability at least $41-\gamma/$ 4, data structure $\mathcal{D}_{0}$ is $\varepsilon/2$-close for the range $[\beta_{\min},\beta_{\circ}]$. In particular, $\hat{Q}(\beta_{\circ}\mid\mathcal{D}_{0})$ is within $e^{\pm\varepsilon/2}$ of $Q(\beta_{\circ})$. Also, by Lemma 17, each $\mathcal{D}_{i}$ is $\varepsilon/2$-close for the range $[\beta_{\circ},\beta_{\max}]$ with probability at least $0.7$. So with probability at least $21-F_{-}(0.7t,0.2t)\geq 1-\gamma/$ 2, at least half the structures $\mathcal{D}_{i}$ are $\varepsilon/2$-close; when this holds, then $\text{median}(\hat{Q}(\alpha\mid\mathcal{D}_{1}),\dots,\hat{Q}(\alpha\mid\mathcal{D}_{t}))$ is within $e^{\pm\varepsilon/2}$ of $\frac{Q(\alpha)}{Q(\beta_{\circ})}$ for all values $\alpha\in[\beta_{\circ},\beta_{\max}]$. Then the final data structure $\mathcal{D}$ is $\varepsilon$-close for the entire range $[\beta_{\min},\beta_{\max}]$. ∎

### v11

@@ 2.1 The Balance subroutine
Given a target $\chi$, we need to find a value $\beta$ with $\mu_{\beta}[0,\chi]\approx 1/2\approx\mu_{\beta}[\chi,n]$. That is, $\chi$ is approximately the median of the distribution $\mu_{\beta}$. We can find the target value $\beta$ via noisy binary search. Formally, for $\tau\in(0,\tfrac{1}{2}),\chi\in[0,n]$, define ${\tt BalancingVals}(\chi,\tau)$ to be the set of values $\beta\in[\beta_{\min},\beta_{\max}]$ satisfying the following two properties:
  • Either $\beta=\beta_{\min}$ or $\mu_{\beta}[0,\chi)\geq\tau$
@@
Note the slight asymmetry with respect to the left and right endpoints. In Appendix D, we show the following result:
Theorem 10.
For $\tau\in(0,\tfrac{1}{2})$ an arbitrary constant, the subroutine ${\tt Balance}(\chi,\tau,\gamma)$ has cost $O(\log\frac{nq}{\gamma})$. With probability at least $1-\gamma$, it returns a value $\beta\in{\tt BalancingVals}(\chi,\tau)$ (we say in this case that the call is good).
Theorem 11.
Let $\tau\in(0,\tfrac{1}{2})$ be an arbitrary constant and let $\varepsilon=(1/2-\tau)/30$. If we are given a data structure $\mathcal{D}$ for ${P^{{\varepsilon}}_{\tt ratio-all}}$, then subroutine ${\tt BalancePreprocessed}(\chi,\tau,\mathcal{D},\gamma)$ has cost $O(\log\frac{q}{\gamma})$. If $\mathcal{D}$ is $\varepsilon$-close, then with probability at least $1-\gamma$, it returns a value $\beta\in{\tt BalancingVals}(\chi,\tau)$.
@@
Consider $\alpha\in[\beta_{\min},\beta_{\max}]$ with $\mu_{\alpha}(x)=\Delta(x)$. The result is clear if $\alpha=\beta$. Suppose that $\alpha<\beta$; the case $\alpha>\beta$ is completely analogous. So $\beta>\beta_{\min}$, and since $\beta\in{\tt BalancingVals}(x,\tau)$, this implies $\mu_{\beta}[0,x)\geq\tau$. We then have:
@@ 2.2 Statistical sampling
For $\varepsilon,\gamma\in(0,\tfrac{1}{2}),\bar{p}\in(0,1]$, suppose we draw random variable $\hat{p}\sim\frac{1}{N}\text{Binom}(N,p)$ where $N\geq\frac{3e^{\varepsilon}\log(4/\gamma)}{(1-e^{-\varepsilon})^{2}\bar{p}}$. Then, with probability at least $1-\gamma$, the following two bounds both hold:
$|\hat{p}-p|\leq\varepsilon(p+\bar{p}),\qquad\text{and}\qquad\hat{p}\in\begin{cases}[e^{-\varepsilon}p,e^{\varepsilon}p]&\mbox{if }p\geq e^{-\varepsilon}\bar{p}\\ [0,\bar{p})&\mbox{if }p<e^{-\varepsilon}\bar{p}\\ \end{cases}$ (2)
@@
Many of our algorithms are based on calling $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta,N)$ and using certain values $\hat{\mu}_{\beta}(I)$ for sets $I\subseteq\mathcal{F}$; they succeed when these estimates $\hat{\mu}_{\beta}(I)$ are close to $\mu_{\beta}(I)$. We say the execution of Sample well-estimates $I$ if Eq. (2) hold for $p=\mu_{\beta}(I)$ and $\hat{p}=\hat{\mu}_{\beta}(I)$; otherwise it mis-estimates $I$. Likewise we say Sample well-estimates $k$ if it well-estimates the singleton set $I=\{k\}$. Since this comes up so frequently, we write
$\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta;\varepsilon,\bar{p},\gamma)\qquad\text{as shorthand for:}\quad\hat{\mu}_{\beta}\leftarrow{\tt Sample}\bigl{(}\beta,\big{\lceil}\tfrac{3e^{\varepsilon}\log(4/\gamma)}{(1-e^{-\varepsilon})^{2}\bar{p}}\big{\rceil}\bigr{)}$
This has cost $O(\frac{\log(1/\gamma)}{\varepsilon^{2}\bar{p}})$, and each set $I$ is well-estimated with probability at least $1-\gamma$. Here $\varepsilon$ should be interpreted as the desired relative error, $\bar{p}$ as the minimum “detection threshold” for $\mu_{\beta}$, and $\gamma$ as the desired failure probability.
In a similar way, our algorithms for ${P^{\delta,\varepsilon}_{\tt count}}$ estimate $\pi(x)$ by sampling some well-chosen value $\mu_{\alpha}(x)$. We record the following straightforward error bound:
@@
Suppose that for $x\in\mathcal{F}$, we have $\alpha\in[\beta_{\min},\beta_{\max}]$ and empirical estimates $\hat{Q}(\alpha),\hat{\mu}_{\alpha}(x)$ satisfying the following bounds:
@@
Suppose we can draw independent samples of non-negative random variables $X_{1},\dots,X_{N}$. Given input $\gamma\in(0,\tfrac{1}{2})$ and $\tau\geq 1$, the procedure ${\tt EstimateProducts}(X,\tau,\gamma)$ uses $O(N\tau\log\tfrac{1}{\gamma})$ total samples of the $X$ variables and returns estimates $(\hat{X}^{\tt prod}_{1},\dots,\hat{X}^{\tt prod}_{N})$. If $\tau\geq\sum_{i=1}^{N}{{\mathbb{S}}[{X_{i}}]}$, then with probability at least $1-\gamma$, it holds that
for all $\text{for all $i=1,\dots,N$: }\frac{\hat{X}_{i}^{\tt prod}}{\prod_{j=1}^{i}{\mathbb{E}}[X_{j}]}\in[e^{-\varepsilon},e^{\varepsilon}]\qquad\qquad\text{for $\varepsilon=\sqrt{\sum_{i=1}^{N}{{\mathbb{S}}[{X_{i}}]}/\tau}$}$
@@ 3 Solving ${P^{{\varepsilon}}_{\tt ratio-all}}$ in the continuous setting
3       draw $X$ from $\mu_{\beta_{i-1}}$ and draw $\eta$ from the Exponential distribution of rate one.
4       set $\beta_{i}=\beta_{i-1}-\eta/X$
5       if $X=0$ or $\beta_{i}<\beta_{\min}$ then return set $B=\{\beta_{1},\dots,\beta_{i-1}\}$
@@
3       draw $\eta$ from the Exponential distribution of rate $k$
4       set $z_{i}=z_{i-1}-\eta$
@@
Let $\mathcal{E}$ denote the event that $\mathcal{D}$ fails to be $\varepsilon$-close. Consider the random process where $B\cap[\beta_{\min},\alpha]$ is revealed while $\alpha$ gradually increases; suppose $\mathcal{E}$ occurs and let $\alpha$ be the first value with $|B(\beta_{\min},\alpha)/k-z(\beta_{\min},\alpha)|>\varepsilon$. Since $z(B)$ is a Poisson process, $B\cap(\alpha,\beta_{\max})$ retains its original, unconditioned probability distribution. Accordingly, we have $|B(\alpha,\beta_{\max})/k-z(\alpha,\beta_{\max})|\leq\varepsilon/2$ with probability at least $21-\gamma/$ 2 , in which case we get
@@ 3.1 Paired Product Estimator
To improve the sample complexity to $O(\frac{q\log n}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$, we need a more advanced statistical technique called the Paired Product Estimator (PPE) developed in [14]. Here, we describe a simple version of the algorithm. [^2] See also [14, 22] for a stand-alone version of the PPE algorithm to directly solve ${P^{{\varepsilon}}_{\tt ratio-point}}$ with high probability.
@@
2 define $\beta_{0}=\beta_{\min}$ and $\beta_{t}=\beta_{\max}$.
@@
4       define random variable $W_{i}=\exp\bigl{(}\frac{\beta_{i}-\beta_{i-1}}{2}\cdot X\bigr{)}$ where $X$ is drawn from $\mu_{\beta_{i-1}}$
5       define random variable $V_{i}=\exp\bigl{(}-\frac{\beta_{i}-\beta_{i-1}}{2}\cdot Y\bigr{)}$ where $Y$ is drawn from $\mu_{\beta_{i}}$
@@
There holds ${\mathbb{E}}[\kappa]\leq\frac{\log(z^{\prime}(\beta_{\max})/z^{\prime}(\beta_{\min}))}{k}.$
@@
As shown in [14, Lemma 3.2], we have $\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\geq\exp(\frac{2\kappa_{i}}{z(\beta_{i-1},\beta_{i})})$ for each $i$ (see also [25]). Equivalently, we have $\kappa_{i}\leq\frac{z(\beta_{i-1},\beta_{i})}{2}\cdot\log\bigl{(}\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\bigr{)}$.
Observe that $z(\beta)$ and $z^{\prime}(\beta)$ are both strictly increasing functions of $\beta$. Thus, for any fixed value $x\in[z^{\prime}(\beta_{\min}),z^{\prime}(\beta_{\max})]$, there is a unique value $\beta\in[\beta_{\min},\beta_{\max}]$ with $z^{\prime}(\beta)=x$, and a unique index $i$ with $z(\beta)\in[z(\beta_{i-1}),z(\beta_{i}))$. Define $A_{x}=z(\beta_{i-1},\beta_{i})$, that is, the distance between the nearest points of the cooling schedule. We can rewrite the bound for $\kappa_{i}$ as
@@
If $k\geq\frac{10\log(z^{\prime}(\beta_{\max})/z^{\prime}(\beta_{\min}))}{\varepsilon^{2}}$ for $\varepsilon\in(0,1)$, then with probability at least $0.7$, the following two bounds hold: (a) $\kappa\leq\varepsilon^{2}$ and (b) $|\log(\hat{W}_{i}^{\tt prod}/\hat{V}_{i}^{\tt prod})-\log Q(\beta_{i})|\leq\varepsilon/2$ for all $i=0,\dots,t$.
@@
$\frac{\hat{V}^{\tt prod}_{i}}{\prod_{j=1}^{i}{\mathbb{E}}[V_{j}]}\in[e^{-\varepsilon/4,\varepsilon/4}],\qquad\frac{\hat{W}^{\tt prod}_{i}}{\prod_{j=1}^{i}{\mathbb{E}}[W_{j}]}\in[e^{-\varepsilon/4,\varepsilon/4}]\qquad\text{for all $i=0,\dots,t$}$
In this case, each value $\hat{Q}(\beta_{i})=\hat{W}^{\tt prod}_{i}/\hat{V}^{\tt prod}_{i}$ is within $[e^{-\varepsilon/2},e^{\varepsilon/2}]$ of the product $\prod_{j=1}^{i}\frac{{\mathbb{E}}[W_{j}]}{{\mathbb{E}}[V_{j}]}$; by telescoping products, this is precisely $Z(\beta_{i})/Z(\beta_{0})=Q(\beta_{i})$. ∎
@@
Consider $\alpha=(1-v)\beta_{i-1}+v\beta_{i}$ for $v\in[0,1]$. Suppose that $v\leq\tfrac{1}{2}$; the case where $v\geq\tfrac{1}{2}$ is completely symmetric. We need to show:
@@ 3.2 A hybrid algorithm
6       call $\mathcal{D}_{0}={\tt TPA-Pratio}(k_{1})$ with $k_{1}=\lceil\frac{40\log 4}{\varepsilon^{2}}\log\tfrac{12}{\gamma}\rceil$ for the interval $[\beta_{\min},\beta_{\circ}]$.
@@
8             Set $\mathcal{D}_{i}\leftarrow{\tt PPE-Pratio}(k_{2})$ with $k_{2}=\lceil\frac{10\log(4n)}{\varepsilon^{2}}\rceil$ for the interval $[\beta_{\circ},\beta_{\max}]$
@@
11output tuple $\mathcal{D}=(\beta_{\circ},\mathcal{D}_{0},\mathcal{D}_{1},\dots,\mathcal{D}_{t})$
    Query algorithm $Q(\alpha\mid{\cal D})$:
@@
By Eq. (6), Theorem 16 and our choice of $k_{1}$, the data structure $\mathcal{D}_{0}$ is $\varepsilon/2$-close for the range $[\beta_{\min},\beta_{\circ}]$ with probability at least $41-\gamma/$ 4. In particular, $\hat{Q}(\beta_{\circ}\mid\mathcal{D}_{0})$ is within $e^{\pm\varepsilon/2}$ of $Q(\beta_{\circ})$. Also, by Lemma 17, each $\mathcal{D}_{i}$ is $\varepsilon/2$-close for the range $[\beta_{\circ},\beta_{\max}]$ with probability at least $0.7$. So with probability at least $21-F_{-}(0.7t,0.2t)\geq 1-\gamma/$ 2, at least half the structures $\mathcal{D}_{i}$ are $\varepsilon/2$-close; when this holds, then $\text{median}(\hat{Q}(\alpha\mid\mathcal{D}_{1}),\dots,\hat{Q}(\alpha\mid\mathcal{D}_{t}))$ is within $e^{\pm\varepsilon/2}$ of $\frac{Q(\alpha)}{Q(\beta_{\circ})}$ for all values $\alpha\in[\beta_{\circ},\beta_{\max}]$. Then the final data structure $\mathcal{D}$ is $\varepsilon$-close for the entire range $[\beta_{\min},\beta_{\max}]$. ∎